# UAV-Assisted Resilience in 6G and Beyond Network Energy Saving: A Multi-Agent DRL Approach

### v1

@@ V-B Coverage Performance Evaluation
Fig. 5 illustrates the total coverage ratio per episode for the proposed MADDPG policy compared against two baseline methods. It can be seen from Fig. 5 that the MADDPG policy (blue line) achieves the highest total coverage ratio in all 15 testing episodes. This superior performance suggests that the agents have successfully learned a complex, cooperative strategy, optimizing their trajectories and power control to serve the maximum number of users in the sleeping cells.
The KNN-Fixed policy (green line), which serves as a stable heuristic, is consistently outperformed by MADDPG. This performance gap can be explained by the policy’s myopic and non-cooperative nature. Lacking coordination, multiple agents can be independently drawn to the same dense user cluster. This fact results in redundant coverage in one area while neglecting users in other areas. The MADDPG agents, in contrast, appear to learn to divide this workload efficiently via the centralized training mechanism.
Conversely, the Random policy (orange line) exhibits extremely poor and highly volatile performance, with its coverage ratio frequently dropping to zero. This result is expected, as its actions have no correlation with the system’s objective. The UAVs are just as likely to fly away from users as they are to fly towards them.
V-C Energy Performance Evaluation
Table I compares the overall energy consumption and user service performance of the proposed MADDPG framework against 3 baseline methods. The total energy consumption is decomposed into the UAV operational energy $E_{\text{UAV}}$ and the energy consumed by active cell $E_{\text{cell}}$.
The All Cell ON baseline represents the scenario in which all GBSs remain fully active at fixed transmission power, consuming the highest amount of energy ($121.06$ Wh). In contrast, other methods that involve UAV assistance or selective GBS activation demonstrate energy savings by dynamically adjusting coverage and transmission behavior. The KNN + Fixed Power and Random baselines consume over $120$ Wh due to their static or non-optimized decision-making, leading to redundant UAV movement or inefficient transmission scheduling. Our MADDPG-based approach achieves the lowest total energy consumption ($91.97$ Wh) while maintaining a comparable user service rate ($78.42\%$). This indicates that the trained agents effectively balance between activating UAVs and GBSs to minimize energy usage without sacrificing network service quality. Specifically, MADDPG reduces the total energy consumption by approximately $24\%$ compared to the All Cell ON configuration, and by around $23.8\%$ compared to the random policy, while serving a higher percentage of users.
It is worth noting that the total served user percentage slightly differs from the coverage rate in the previous section. Here, the metric accounts for all users successfully served by either UAVs or active GBSs, whereas the coverage rate previously measured only the UAV-needed coverage proportion. Overall, the MADDPG-based coordination strategy achieves the best trade-off between energy efficiency and service performance, confirming its capability to dynamically adapt to traffic variations and power constraints in green UAV-assisted cellular networks.
@@ VI Conclusion
This paper presents a UAV-assisted approach to green, resilient wireless networks using a MADDPG framework with CTDE. The method jointly optimizes UAV trajectories, transmit power, and user-UAV association under a sleeping-cell strategy to maximize user coverage while minimizing UAV energy; in simulations it outperforms conventional baselines, achieving the highest coverage and reducing total energy consumption by 24% versus an All-Cell-ON configuration. Future work includes lightweight policy distillation and transfer learning to speed convergence; integrating adaptive rate control, user prioritization, and dynamic bandwidth allocation to better satisfy per-user QoS; and jointly optimizing the GBS sleep schedule with UAV trajectories and transmit power. Overall, integrating MADRL with UAV-assisted architectures is a promising path toward sustainable, autonomous, and resilient 6G-era networks.

### v2

@@ V-C Energy Performance Evaluation
A striking observation is that the QoS-Aware Greedy baseline consumes the highest total energy ($126.65$ Wh), even exceeding the All Cell ON case ($121.06$ Wh). This is due to the UAVs’ excessive maneuvering ($E_{\text{UAV}}\approx 40.28$ Wh) to chase optimal SINR positions without regard for propulsion costs. This highlights a critical insight: employing UAVs without intelligent energy-aware control can be detrimental to network sustainability.
In strong contrast, our MADDPG approach achieves the lowest total energy ($91.97$ Wh), reducing consumption by approximately 27.38% compared to Greedy and 24.03% compared to the All Cell ON baseline. By learning to position UAVs strategically rather than aggressively, MADDPG delivers virtually the same coverage as the Greedy approach ($78.37\%$ vs $78.42\%$) but with $7x$ less UAV flight energy ($5.60$ Wh vs $40.28$ Wh). This confirms that the proposed MADDPG framework achieves the best trade-off between energy efficiency and service performance, making it a viable strategy for enabling resilient, energy-efficient 6G networks.
@@ VI Conclusion
This paper presents a UAV-assisted resilience framework for 6G energy-saving networks using a CTDE-based MADDPG approach. The proposed method enables UAVs to autonomously adapt their trajectories, transmit power, and user association to dynamically compensate for coverage holes induced by traffic-aware GBS sleeping strategies. Simulation results demonstrate that our framework achieves a superior trade-off between resilience and sustainability, attaining user coverage rates comparable to an aggressive QoS-Aware Greedy baseline while consuming drastically less energy, saving approximately 24% compared to the conventional All-Cell-ON configuration. These findings highlight that while heuristic greedy strategies can effectively restore coverage, they are energetically unsustainable for UAV-assisted networks due to excessive propulsion costs. Future work includes lightweight policy distillation to speed convergence, integrating dynamic bandwidth allocation to better satisfy per-user QoS, and jointly optimizing GBS sleep schedules. Overall, this study confirms that integrating MADRL with UAV architectures is a key enabler for autonomous, green, and resilient 6G systems.